4. If an individual’s preferences are described by the utility function U(X1 , X2 ) = X12 + X22,

a) graph the indifference curve for U = 20 and U= 40.

b) Find the optimal consumption quantities if P1 = US$2.50 ; P2 = US$ 7.50; and M = US$ 60.



1
Expert's answer
2021-06-21T11:58:47-0400

a)

"U(X_1," "X_2)\ =\ X_1^2+X_2^2"


"U=20"


"20=X_1^2+X_2^2"


"X_1^2=20-X_2^2"


For U=40


"40=X_1^2+X_2^2"


"X_1^2=40-X_2^2"




b)

Budget line equation= "P_1X_1\ +P_2X_2=M" .................Equation 1


Substituting the values given to Equation 1 above. We get,


"2.50X_1+7.50X_2=60..........................2"

"U(X_1,X_2)=X_1^2+X_2^2"



"MRS(X_1,X_2)=\frac{MUx_!}{MUx_2}"


"MUX_1=X_1,MUX_2" "=2X_2"


"MRS=X_1,X_2=\frac{X_1}{X_2}"


Optimal Bundle=


"\frac{P_1}{P_2}=MRS""(X_1,X_2)"


"\frac{2.5}{7.5}=\frac{X_1}{X_2}"


"2.5X_2=7.5X_1"


"X_2=3X_1.......................................3"


Substituting "X_2" in equation 2


We get,


"X_1=\frac{60}{25}"


"=2.4"

Substituting "X_1" in equation 3


We get,


"X_2=3X_1"


"=3\times2.4=7.2"


"X_1=2.4,\ X_2=7.2"